OpenStudy (jkbo):

What is the geometric mean of 7 and 12? (1 point) a.1 b.9.5 c.2√21 d.4√21

OpenStudy (anonymous):

Ex: If (a,b,c) is a geometric sequence then b is a G.M b^2=a*c

OpenStudy (anonymous):

can u get it now ?

OpenStudy (jkbo):

no i still dont gett it

OpenStudy (anonymous):

According to the G.M of a G.S: let b is the G.m between 7,12 b^2=a*c b^2=7*12 b^2=84 b=84^1/2= c.2√21

OpenStudy (jkbo):

thank u